7Find the energy of the photons for each of the fol
Answer #1

i)

Wave length of radio wave =\lambda =80cm=0.8m

Energy of photon =E

We know

E=\frac{hc}{\lambda }

h= Planck's constant =6.626\times 10^{-34}J.s

c= Speed of light =3\times 10^{8}m/s

E=\frac{6.626\times 10^{-34}\times 3\times 10^{8}}{0.8}J=24.85\times 10^{-26}J

ii)

Wave length of violet light =\lambda =390nm=390\times 10^{-9}m

Energy of photon =E

We know

E=\frac{hc}{\lambda }

h= Planck's constant =6.626\times 10^{-34}J.s

c= Speed of light =3\times 10^{8}m/s

E=\frac{6.626\times 10^{-34}\times 3\times 10^{8}}{390\times 10^{-9}}J=50.97\times 10^{-19}J

iii)

Wave length of X-ray =\lambda =0.25nm=0.25\times 10^{-9}m

Energy of photon =E

We know

E=\frac{hc}{\lambda }

h= Planck's constant =6.626\times 10^{-34}J.s

c= Speed of light =3\times 10^{8}m/s

E=\frac{6.626\times 10^{-34}\times 3\times 10^{8}}{0.25\times 10^{-9}}J=79.51\times 10^{-17}J